OpenStudy (anonymous):

2. A chemist wishes to make 3 liters of a 7% acid solution by mixing a 9% acid solution and a 4% acid solution. How many liters of each solution should the chemist use? Write your answer as a complete sentence. Be sure to: • Define your variable and expressions for the quantities. • Write an equation that models the problem. • Solve the equation. • State the answer in a complete sentence. Answer:

OpenStudy (anonymous):

i will help, but you have to define the variable

OpenStudy (anonymous):

okay you want me to go first or you?

OpenStudy (anonymous):

you define the variable name it, and say what it represents, we can take it from there

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

variables is y=unknown

OpenStudy (anonymous):

?

OpenStudy (anonymous):

y equals what? amount of 9% solution, or the amount of 4% solution?

OpenStudy (anonymous):

i would say 4% solution?

OpenStudy (anonymous):

so for Define your variable and expressions for the quantities. -4% solution

OpenStudy (anonymous):

damn i read it wrong since the total is 3, the amount of 9% solution is \(3-y\)

OpenStudy (anonymous):

lol thats ok lets keep going if you want

OpenStudy (anonymous):

you will have 4% of y and 9% of \(3-y\) for a total of \[.04y+.09(3-y)\] which you want to be equal to 7% of 3 set \[.04y+.09(3-y)=.07\times 3\]

OpenStudy (anonymous):

easier if you remove decimals by multiplying by 100 and start with \[4y+9(3-y)=7\times 3\] solve for \(y\)

OpenStudy (anonymous):

Define your variable and expressions for the quantities. -4y+9(3−y)=7×3